OpenStudy (shaik0124):

A pump can fill a tank with water in 2 hours. Because of a leak, it took 2 hours to fill the tank. The leak can drain all the water of the tank in: A. 4 1/ 3 hours B. 7 hours C. 8 hours D. 14 hours

OpenStudy (shaik0124):

Can u please help

OpenStudy (dtan5457):

The leak took the same amount of time?

OpenStudy (mathstudent55):

Reread the problem. It makes no sense.

OpenStudy (mathstudent55):

I guess the leak was already fixed. I don't know about math, but the plumber is good.

OpenStudy (anonymous):

@mathstudent55 LOL

OpenStudy (shaik0124):

A pump can fill a tank with water in 2 hours. Because of a leak, it took 2 1/3 hours to fill the tank. The leak can drain all the water of the tank in:

OpenStudy (shaik0124):

@mathstudent55

OpenStudy (shaik0124):



OpenStudy (mathstudent55):

The pump's speed is 1 tank in 2 hours or 0.5 tank/hour. With the leak, the speed of filling the tank is 1 tank/( 2 1/3 hours) = 3/7 tank/hour. The difference is 1/2 - 3/7 = 4/14 = 1/7 tank per hour. The leak is 1/7 tank per hour. That means it will take 7 hours to for the leak to empty the tank.

OpenStudy (shaik0124):

1/2 - 3/7 it's1/14 check once

OpenStudy (shaik0124):

y u r taking the difference ??@mathstudent55

OpenStudy (mathstudent55):

Wow. You are correct. I subtracted incorrectly and reduced incorrectly. Let me do it again. BTW, yes, I am subtracting.

OpenStudy (shaik0124):

It's ok math .reason y u r taking the difference of two

OpenStudy (shaik0124):

1/2 -3/7 ????

OpenStudy (mathstudent55):

The pump's speed is 1 tank in 2 hours or 1/2 tank/hour. With the leak, the speed of filling the tank is 1 tank/( 2 1/3 hours) = 3/7 tank/hour. The difference (subtraction) is 1/2 - 3/7 = 7/14 - 6/14 = 1/14 tank per hour. The leak is 1/14 tank per hour. That means it will take 14 hours to for the leak to empty the tank.

OpenStudy (mathstudent55):

Thanks for pointing that out. See, I did say above I didn't know about the math, but the plumber was good.